Please help me write a function for the points (1,1) and (3,-1)
Answer by stanbon(75887)

(1,1) and (3,-1)
slope = (-1-1)/(3-1)= -1
Use this slope and a point to find "b"
of y=mx+b
1=(-1)1+b
b=2
EQUATION: y=-x+2
Cheers,
Stan H.